OpenStudy (anonymous):

13^2= square root or 177+b^2?????

OpenStudy (anonymous):

okey dokey

OpenStudy (anonymous):

169 = 177+b^2 subtract 177 from both sides to isolate b^2

OpenStudy (anonymous):

oops, sorry. I read your post wrong

OpenStudy (anonymous):

\[169 = \sqrt{177+b ^{2}}\]

OpenStudy (anonymous):

square both sides to find \[28561=177+b ^{2}\]

OpenStudy (anonymous):

subtract 177 from both sides: 28394=b^2

OpenStudy (anonymous):

take the square root of both sides to solve for b \[\sqrt{28384} = b = 168.476\] (approximately)

OpenStudy (anonymous):

oh so i don't square root the 177 til the end???

OpenStudy (anonymous):

was the original problem \[13^{2}=\sqrt{177+b ^{2}}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

We never had to find the square root of 177 by itself

OpenStudy (anonymous):

If you square a square root, the square root goes away. That's what I did to find 28561=177+b^2

OpenStudy (anonymous):

from there I subtracted 177 from both sides to get the variable by itself

OpenStudy (anonymous):

so then the equation becomes \[28384 = b ^{2}\]

OpenStudy (anonymous):

then to solve for b, i took the square root of both sides